# Improved Extended Regular Expression Matching

### v1

@@ 5 Reducing Space
We now show how to reduce the $O(n^{2}k/w+n+m)$ space of the algorithm from the previous section to $O(n^{2}\log k/w+n+m)$. We modify the bottom-up traversal to instead use a top-down recursive traversal, where we always process the child cluster with the largest subtree in the macro tree and carefully discard match graphs during the algorithm.
@@ 5.1 Heavy Path Decomposition
Let $M$ be the macro tree for $R$. Given a cluster $C\in CS$, define the size of $C$ to be the number of descendant clusters of $C$ in $M$. As in the heavy-path decomposition of Sleator and Tarjan [46], we classify each cluster in $M$ as either heavy or light. The root of $M$ is light, and for each internal cluster $C$ with two children, we pick a child cluster of maximum size and classify it as heavy. The remaining children are light. An edge to a light child is a light edge and an edge to a heavy child is a heavy edge. Removing the light edges, we partition $M$ into heavy paths.
@@ 5.2 Macro Tree Traversal
Since processing a cluster only requires the match graphs of children of a cluster, it follows that the algorithm correctly solves the problem as before.
We show that this reduces the number of stored match graphs to $O(\log k)$. Consider the algorithm at the time of processing a cluster $C$, and let $L$ be the path of clusters from the root of the macro tree to $C$. By the ordering, the algorithm stores at this point the match graphs of the children of $C$ and for each light cluster on $L$, the match graph of its heavy sibling. By Lemma 7, we thus store at most $\log\ell+O(1)=O(\log k)$ match graphs. In total, the algorithm uses $O(n^{2}\log k/w+n+m)$ space.
@@ 6 Black-Box TNFA Simulation
Nearly all of the existing (non-extended) regular expression matching algorithms are TNFA simulation algorithms [39, 8, 7, 13]. The state-of-the-art algorithms achieve either $t(n,m)=O(nm/\log n+n+m)$ and $s(n,m)=O(m+n^{\varepsilon})$, for constant $\varepsilon$, $0<\varepsilon<1$ [8] or $t(n,m)=O(nm\log w/w+n+m)$ and $s(n,m)=O(m)$ [7, 13]. Note that the former result is slightly different than stated in the original paper, see [9] for how to improve this. Plugging these into our algorithm, we improve the $n^{2}m$ term in our time bound to $O(n^{2}m/\min(w/\log w,\log n))$. Combined with our improved space bound from Section 5, this completes the proof of Theorem 1.

### v2

@@ 5 Reducing Space
We now show how to reduce the $O(n^{2}k/w+n+m)$ space of the algorithm from the previous section to $O(n^{2}\log k/w+n+m)$. We modify the bottom-up traversal to use a top-down recursive traversal, where we always process the child cluster with the largest subtree in the macro tree and carefully discard match graphs as the algorithm proceeds.
@@ 5.1 Heavy Path Decomposition
Let $M$ be the macro tree for $R$. Given a cluster $C\in CS$, define the weight of $C$ to be the number of descendant clusters of $C$ in $M$. As in the heavy-path decomposition of Sleator and Tarjan [55], we classify each cluster in $M$ as either heavy or light. The root of $M$ is light, and for each internal cluster $C$ with two children, we pick a child cluster of maximum weight and classify it as heavy. The remaining children are light. An edge to a light child is a light edge and an edge to a heavy child is a heavy edge. Removing the light edges, we partition $M$ into heavy paths.
@@ 5.2 Macro Tree Traversal
Since processing a cluster requires only the match graphs of its children, the algorithm correctly solves the problem as before. We show that this reduces the number of stored match graphs to $O(\log k)$. Consider the algorithm at the time of processing a cluster $C$, and let $L$ be the path of clusters from the root of the macro tree to $C$. By the ordering, the algorithm stores, at this point, the match graphs of the children of $C$ and, for each light cluster on $L$, the match graph of its heavy sibling. Recall from Lemma 2 that $M$ contains $\ell=O(k)$ nodes. Hence, by Lemma 7, we thus store at most $\log\ell+O(1)=O(\log k)$ match graphs. In total, the algorithm uses $O(n^{2}\log k/w+n+m)$ space.
@@ 6 Black-Box TNFA Simulation
Nearly all of the existing (non-extended) regular expression matching algorithms are TNFA simulation algorithms [46, 9, 8, 14]. The state-of-the-art algorithms achieve either $t(n,m)=O(nm/\log n+n+m)$ and $s(n,m)=O(m+n^{\varepsilon})$, for constant $\varepsilon$, $0<\varepsilon<1$ [9] or $t(n,m)=O(nm\log w/w+n+m)$ and $s(n,m)=O(m)$ [8, 14]. Note that the former result is slightly different than stated in the original paper, see [10] for how to improve this. Plugging these into our algorithm, we improve the $n^{2}m$ term in our time bound to $O(n^{2}m/\max(w/\log w,\log n))$. Combined with our improved space bound from Section 5, this completes the proof of Theorem 1.
7 Interval Operator
We now show how to extend our solution to efficiently support the interval operator. We do this by adding the operator to our match graphs and plugging it into our algorithm. Our solution is surprisingly simple and demonstrates the power of our framework.
Let $R$ be a regular expression. For integers $a$ and $b$, $0\leq a\leq b$, the interval operator, $R^{[a,b]}$ specifies the set of strings that consists of $\ell$ concatenations of $R$, where $\ell\in[a,b]$. Thus, the language $L(R^{[a,b]})$ is given by $L(R^{[a,b]})=\bigcup_{a\leq i\leq b}L(R)^{i}$ (recall that $L(R)^{i}$ is the set of strings from $R$ concatenated with itself $i$ times). As a shorthand, we write $R^{a}$ to denote $R^{[a,a]}$. We have the following properties of the interval operator that follow immediately from the definition.
Lemma 9.
Let $R$ be a regular expression and let $a$ and $b$ be integers such that $0\leq a\leq b$. Then,
  • (i)
    $R^{[a,b]}=R^{a}\odot R^{[0,b-a]}$,
  • (ii)
    $R^{a}\odot R^{b}=R^{a+b}$, and
  • (iii)
    $R^{[0,a]}\odot R^{[0,b]}=R^{[0,a+b]}$.
We extend our match graphs from Section 4.2 to include the interval operator. Let $G$ be a match graph on vertices $g_{1},\ldots,g_{n+1}$. For the match graph $G^{[a,b]}$, we add an edge $(g_{i},g_{j})$ for each pair $i,j$ with a directed path from $g_{i}$ to $g_{j}$ in $G$ with length $\ell$, where $a\leq\ell\leq b$. Using the properties of Lemma 9 and repeated doubling, we show how to compute $G^{[a,b]}$ from $G$ with $O(\log b)$ matrix multiplications.
Lemma 10.
Let $G$ be a match graph on a string of length $n$ and let $a$ and $b$ be integers such that $0\leq a\leq b$. Then, we can store $G$ in $O(n^{2}/w+1)$ space and compute $G^{[a,b]}$ in $O(n^{\omega}\log b)$ time, for any integers $a$ and $b$, $0\leq a\leq b$.
Proof.
Let $G$, $a$, and $b$ be given as above. As in Section 3, we store match graphs as boolean matrices in $O(n^{2}/w+1)$ space. Let $j$ be the largest integer such that $2^{j}\leq a$. We first compute $G^{a}$. To do so, we compute the sequence of match graphs $G^{1}$, $G^{2}$, …, $G^{2^{j}}$ from $G$ as follows. We have $G^{1}=G$ and for $i=1,\ldots j$, we compute
$G^{2^{i}}=G^{2^{i-1}}\odot G^{2^{i-1}}.$ (1)
Now, let $r_{1},\ldots,r_{z}$ be the sequence of $1$ bits in the binary representation of $a$, and compute
$G^{a}=\bigodot_{1\leq i\leq z}G^{2^{r_{i}}}.$ (2)
Correctness for (1) and (2) follows from Lemma 9(ii).
We compute $G^{[0,b-a]}$ using the same approach. Let $j^{\prime}$ be largest integer such that $2^{j^{\prime}}\leq b-a$. We compute the sequence of match graphs $G^{[0,1]}$, $G^{[0,2]}$, …, $G^{[0,2^{j^{\prime}}]}$, where we construct $G^{[0,1]}$ from $G$ by adding self-loops to each node, and $G^{[0,2^{i}]}$, for $i=1,\ldots,j^{\prime}$, as $G^{[0,2^{i}]}=G^{[0,2^{i-1}]}\odot G^{[0,2^{i-1}]}$ using Lemma 9(iii). We then compute $G^{[0,b-a]}$ by concatenating the match graphs of the match graph sequence corresponding to the binary representation of bab-a as above. Finally, we compute $G^{[a,b]}=G^{a}\odot G^{[0,b-a]}$ according to Lemma 9(i).
To compute $G^{a}$ and $G^{[0,b-a]}$ we use $O(\log a+\log(b-a))=O(\log b)$ match graph concatenations. The remaining step of the algorithm uses a constant number of operations. By Lemma 3, the total time is $O(n^{\omega}\log b)$. We can avoid storing the $O(j)$ match graphs during the computation of $G^{a}$ by computing (2) in parallel with $G^{1}$, $G^{2}$, …, $G^{2^{j}}$ in left-to-right order and only storing a constant number of previously computed match graphs. The same applies when computing $G^{[0,b-a]}$. Thus, the total space is $O(n^{2}/w+1)$. ∎
We plug the match graph construction from Lemma 10 for the interval operator into our main algorithm of Section 4.3 whenever we compute $G(p)$ in step 1. All of the other steps remain the same. The further techniques in Sections 5 and 6 still apply. In summary, we have shown the following result.
Theorem 11.
Given an extended regular expression $R$ of length $m$ containing $k$ extended operators, $k^{\prime}$ interval operators, and a string $Q$ of length $n$, we can solve the extended regular expression matching problem for $R$ and $Q$ in space $O(\frac{n^{2}\log k}{w}+n+m)=O(n^{2}+m)$ and time
$O\left(n^{\omega}(k+k^{\prime}\log B)+\frac{n^{2}m}{\max(w/\log w,\log n)}+m\right).$
Here, $B$ is the maximum of the upper bounds in ranges for the $k^{\prime}$ interval operators in $R$.
Thus, compared to Theorem 1, we only incur an additional $\log B$ factor on each of the $k^{\prime}$ interval operators.